Find f(0),f(2),f(-2),f(3),f(√2 ), and f(3t) if :

a) f(x)=3x^2 -2

b) f(x)=√(x+1)


1
Expert's answer
2020-05-11T10:59:06-0400


a) "f(x)=3x^2-2,"

"f(0)=3*0^2-2=-2"

"f(2)=3*2^2-2=3*4-2=10"

"f(-2)=3*(-2)^2-2=3*4-2=10"

"f(3)=3*3^2-2=3*9-2=25"

"f(\sqrt{2})=3*(\sqrt{2})^2-2=3*2-2=4"

"f(3t)=3*(3t)^2-2=3*9t^2-2=27t^2-2"


b) "f(x)=\sqrt{x+1},"

"f(0)=\sqrt{0+1}=1"

"f(2)=\sqrt{2+1}=\sqrt{3}"

"f(-2)=\sqrt{-2+1}=\sqrt{-1}" is undetermined in real numbers

"f(3)=\sqrt{3+1}=\sqrt{4}=2"

"f(\sqrt{2})=\sqrt{\sqrt{2}+1}"

"f(3t)=\sqrt{3t+1}"